### v4

@@ Appendix A Auxiliary Results
On the one hand, via (6), (45) and Lemma A.1(i), using that $(2+\delta)V_{0}^{\prime\prime}(\xi^{0}_{\star})\gamma_{1}>1+\frac{\delta}{2}$ if $\beta=1$ by (46),
@@
$S^{2,2}_{h}$ $\coloneqq\sum_{k=1}^{\lceil h^{-2}\rceil}\mathbb{E}\bigg[\Big(\frac{h^{-1}}{\lceil h^{-2}\rceil}\eta^{h}_{k}\Big)^{2}\bigg|\mathcal{F}^{h}_{k-1}\bigg]=\frac{h^{-2}}{\lceil h^{-2}\rceil}\frac{\mathbb{V}\mathrm{ar}\big((X_{h}-\xi^{h}_{\star})^{+}\big)}{(1-\alpha)^{2}},$
@@
$\Sigma_{n}=\frac{1}{\gamma_{n}}\sum_{k=1}^{n}\gamma_{k}^{2}\Pi_{k+1:n}^{2}.$ (52)
@@
$\lim_{\mathcal{H}\ni h\downarrow 0}S^{1,1}_{h}=\frac{\gamma_{1}\alpha}{1-\alpha}\lim_{n\uparrow\infty}\Sigma_{n},$ (53)
@@
$\frac{\gamma_{n}-\gamma_{n+1}}{\gamma_{n+1}}=\frac{\mathds{1}_{\beta=1}}{\gamma_{1}}\gamma_{n}+\mathrm{o}(\gamma_{n}),\quad\gamma_{n+1}-\gamma_{n}=\mathrm{o}(\gamma_{n})\quad\mbox{ and }\quad\gamma_{n}\gamma_{n+1}=\mathrm{o}(\gamma_{n}).$
@@
$\Sigma_{n+1}=\Sigma_{n}+\Big(1-\Big(2V_{0}^{\prime\prime}(\xi^{0}_{\star})-\frac{\mathds{1}_{\beta=1}}{\gamma_{1}}\Big)\Sigma_{n}\Big)\gamma_{n}+(\Sigma_{n}+1)\mathrm{o}(\gamma_{n}).$
@@
$\Sigma_{\star}\coloneqq\frac{1}{2V_{0}^{\prime\prime}(\xi^{0}_{\star})-\frac{\mathds{1}_{\beta=1}}{\gamma_{1}}}.$ (54)
@@
$\Delta\Sigma_{n}\coloneqq\Sigma_{n}-\Sigma_{\star},$ (55)
@@
$=\big(1-\mu\gamma_{n}+\mathrm{o}(\gamma_{n})\big)\Delta\Sigma_{n}+\mathrm{o}(\gamma_{n}),$
@@
$|\Delta\Sigma_{n+1}|\leq\big(1-(\mu+\varepsilon)\gamma_{n}\big)|\Delta\Sigma_{n}|+\varepsilon\gamma_{n}.$ (56)
@@
$\lim_{n\uparrow\infty}\Sigma_{n}=\Sigma_{\star}.$ (57)
Combining (53) and (57),
$S^{1,1}_{h}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathbb{P}\text{\rm-as}}}}\frac{\alpha}{1-\alpha}\,\frac{\gamma_{1}}{2V_{0}^{\prime\prime}(\xi^{0}_{\star})-\frac{\mathds{1}_{\beta=1}}{\gamma_{1}}}\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
Given that $x\mapstochar\rightarrow x^{+}$ is $1$-Lipschitz on $\mathbb{R}$,
@@
$\mathbb{E}[\mathbb{V}\mathrm{ar}(\varphi(Y,Z)|Y)]\leq\mathbb{E}[\mathbb{V}\mathrm{ar}(\varphi(Y,Z)|Y)]+\mathbb{V}\mathrm{ar}(\mathbb{E}[\varphi(Y,Z)|Y])=\mathbb{V}\mathrm{ar}(\varphi(Y,Z))<\infty.$
@@
$\mathbb{E}[(X_{h}-X_{0})^{2}]=h\,\mathbb{E}[\mathbb{V}\mathrm{ar}(\varphi(Y,Z)|Y)]\rightarrow 0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$ (58)
@@
$\mathbb{V}\mathrm{ar}\big((X_{h}-\xi^{h}_{\star})^{+}\big)\rightarrow\mathbb{V}\mathrm{ar}\big((X_{0}-\xi^{0}_{\star})^{+}\big)\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
$S^{2,2}_{h}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathbb{P}\text{\rm-as}}}}\frac{\mathbb{V}\mathrm{ar}\big((X_{0}-\xi^{0}_{\star})^{+}\big)}{(1-\alpha)^{2}}\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
$\mathbb{E}\big[e^{h}_{k}\eta^{h}_{k}\big|\mathcal{F}^{h}_{k-1}\big]=\frac{\alpha}{(1-\alpha)^{2}}\,\mathbb{E}[(X_{h}-\xi^{h}_{\star})^{+}]\rightarrow\frac{\alpha}{(1-\alpha)^{2}}\,\mathbb{E}[(X_{0}-\xi^{0}_{\star})^{+}]\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$ (59)
@@
$\lim_{\mathcal{H}\ni h\downarrow 0}\frac{h^{-(1+\beta)}}{\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}\gamma_{k}\Pi_{k+1:\lceil h^{-2}\rceil}\mathbb{E}\big[e^{h}_{k}\eta^{h}_{k}\big|\mathcal{F}^{h}_{k-1}\big]=\frac{\alpha\,\mathbb{E}[(X_{0}-\xi^{0}_{\star})^{+}]\mathds{1}_{\beta=1}}{(1-\alpha)^{2}}\lim_{n\uparrow\infty}\widetilde{\Sigma}_{n},$ (60)
@@
$\lim_{n\uparrow\infty}{\sum_{k=1}^{n}\gamma_{k}\Pi_{k+1:n}}=\widetilde{\Sigma}_{\star}$. (61)
Finally, by (60) and (61),
$S^{1,2}_{h}=S^{2,1}_{h}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathbb{P}\text{\rm-as}}}}\frac{\alpha\,\mathbb{E}[(X_{0}-\xi^{0}_{\star})^{+}]\mathds{1}_{\beta=1}}{V_{0}^{\prime\prime}(\xi^{0}_{\star})(1-\alpha)^{2}}\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
By the decomposition (35),
$\xi^{h}_{n}-\xi^{h}_{\star}=\frac{1}{V_{h}^{\prime\prime}(\xi^{h}_{\star})}a^{h}_{n}+\frac{1}{V_{h}^{\prime\prime}(\xi^{h}_{\star})}r^{h}_{n}+\frac{1}{V_{h}^{\prime\prime}(\xi^{h}_{\star})}\rho^{h}_{n}+\frac{1}{V_{h}^{\prime\prime}(\xi^{h}_{\star})}e^{h}_{n},$ (62)
recalling the definitions (37)–(39), with $\{a^{h}_{n},n\geq 1\}$ given by
$a^{h}_{n}=-\frac{1}{\gamma_{n}}\big(\xi^{h}_{n}-\xi^{h}_{\star}-(\xi^{h}_{n-1}-\xi^{h}_{\star})\big).$ (63)
From (13) and (62),
@@
$\frac{1}{n}\sum_{k=1}^{n}\frac{1}{\gamma_{k}}\big(\xi^{h}_{k}-\xi^{h}_{\star}-(\xi^{h}_{k-1}-\xi^{h}_{\star})\big)$ $=\frac{1}{n}\Big(\frac{1}{\gamma_{n}}(\xi^{h}_{n}-\xi^{h}_{\star})-\frac{1}{\gamma_{1}}(\xi^{h}_{0}-\xi^{h}_{\star})\Big)$ (64)
@@
$\mathbb{E}\bigg[\bigg|\frac{1}{n}\sum_{k=2}^{n}\Big(\frac{1}{\gamma_{k-1}}$ $-\frac{1}{\gamma_{k}}\Big)(\xi^{h}_{k-1}-\xi^{h}_{\star})\bigg|\bigg]$ (65)
@@
$\mathbb{E}\bigg[\bigg|\frac{1}{n}\sum_{k=1}^{n}\frac{1}{\gamma_{k}}a^{h}_{k}\bigg|\bigg]\leq\frac{C}{n\sqrt{\gamma_{n}}}.$ (66)
@@
$\frac{h^{-1}}{V_{h}^{\prime\prime}(\xi^{h}_{\star})\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}a^{h}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
$\frac{h^{-1}}{V_{h}^{\prime\prime}(\xi^{h}_{\star})\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}r^{h}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
Recalling that $\{\rho^{h}_{k},k\geq 1\}$ is a sequence of martingale increments, following (49),
@@
$\frac{h^{-1}}{V_{h}^{\prime\prime}(\xi^{h}_{\star})\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}\rho^{h}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
We start by checking the conditional Lindeberg condition. Note that Step 7.1 of the proof of Theorem 1.2, Appendix B, already guarantees that $\big\{\frac{h^{-1}}{\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}\eta^{h}_{k},h\in\mathcal{H}\big\}$ satisfies this condition. Moreover, by assumption, via (39) and (6),
@@
$\leq\frac{c_{\alpha}^{2+\delta}}{|V_{h}^{\prime\prime}(\xi^{h}_{\star})|^{2+\delta}\lceil h^{-2}\rceil^{\frac{\delta}{2}}}\rightarrow 0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0,$
@@
$S^{2,2}_{h}$ $\coloneqq\sum_{k=1}^{\lceil h^{-2}\rceil}\mathbb{E}\bigg[\Big(\frac{h^{-1}}{\lceil h^{-2}\rceil}\eta^{h}_{k}\Big)^{2}\bigg|\mathcal{F}^{h}_{k-1}\bigg]=\frac{h^{-2}}{\lceil h^{-2}\rceil}\frac{\mathbb{V}\mathrm{ar}\big((X_{h}-\xi^{h}_{\star})^{+}\big)}{(1-\alpha)^{2}},$
@@
$S^{1,1}_{h}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathbb{P}\text{\rm-as}}}}\frac{\alpha(1-\alpha)}{f_{X_{0}}(\xi^{0}_{\star})^{2}}\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
Finally, using that $\lim_{\mathcal{H}\ni h\downarrow 0}V_{h}^{\prime\prime}(\xi^{h}_{\star})=V_{0}^{\prime\prime}(\xi^{0}_{\star})=(1-\alpha)^{-1}f_{X_{0}}(\xi^{0}_{\star})$ and (59), we obtain by Cesàro’s lemma that
$S^{1,2}_{h}=S^{2,1}_{h}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathbb{P}\text{\rm-as}}}}\frac{\alpha}{1-\alpha}\frac{\mathbb{E}[(X_{0}-\xi^{0}_{\star})^{+}]}{f_{X_{0}}(\xi^{0}_{\star})}\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
$U_{\ell}\coloneqq\begin{pmatrix}\frac{1}{(KM^{\ell-1}(M-1))^{\frac{1}{2}}}\sum_{k=KM^{\ell-1}+1}^{KM^{\ell}}\Phi^{(k)}(Y)\\ \frac{1}{(KM^{\ell-1})^{\frac{1}{2}}}\sum_{k=1}^{KM^{\ell-1}}\Phi^{(k)}(Y)\end{pmatrix}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathcal{L}}}}\mathcal{N}\big(0,\mathbb{E}[\Phi^{(1)}(Y)^{2}|Y]I_{2}\big),\quad\mbox{ as }\quad\ell\uparrow\infty,$
@@
$h_{\ell}^{-\frac{1}{2}}(X_{h_{\ell}}-X_{h_{\ell-1}})\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathcal{L}}}}\mathcal{N}\big(0,(M-1)\mathbb{V}\mathrm{ar}(\varphi(Y,Z)|Y)\big)\quad\mbox{ as }\quad\ell\uparrow\infty,$ (67)
@@
$\mathbb{E}[\mathbb{V}\mathrm{ar}(\varphi(Y,Z)|Y)]\leq\mathbb{E}[\mathbb{V}\mathrm{ar}(\varphi(Y,Z)|Y)]+\mathbb{V}\mathrm{ar}(\mathbb{E}[\varphi(Y,Z)|Y])=\mathbb{V}\mathrm{ar}(\varphi(Y,Z))<\infty,$ (68)
so that $\mathbb{V}\mathrm{ar}(\varphi(Y,Z)|Y)<\infty$ -$\mathbb{P}\text{\rm-as}$. This concludes the proof.
@@
$\mathbb{P}(X_{h_{\ell-1}}\leq\xi<X_{h_{\ell}})$ $=\mathbb{P}\big(X_{h_{\ell-1}}\leq\xi<X_{h_{\ell-1}}+h_{\ell}^{\frac{1}{2}}G_{\ell}\big)$ (69)
@@
$\mathbb{P}(X_{h_{\ell}}\leq\xi<X_{h_{\ell-1}})=h_{\ell}^{\frac{1}{2}}\mathbb{E}[G_{\ell}^{-}f_{X_{h_{\ell-1}}|G_{\ell}}(\xi)]+h_{\ell}^{\frac{1}{2}}r_{\ell}^{-},$ (70)
@@
$\mathbb{E}[|G_{\ell}|f_{X_{h_{\ell-1}}|G_{\ell}}(\xi)]$ $=\mathbb{E}[|G_{\ell}|(f_{X_{h_{\ell-1}}|G_{\ell}}(\xi)-f_{G_{\ell}}(\xi))\mathds{1}_{|G_{\ell}|\leq C}]$ (71)
@@
Using that $\big\{g\mapstochar\rightarrow f_{X_{h_{\ell-1}}|G_{\ell}=g}(\xi),\ell\geq 1\big\}$ are bounded uniformly in $\ell\geq 1$ and converge locally uniformly towards $g\mapstochar\rightarrow f_{g}(\xi)$, by the dominated convergence theorem,

### v5

@@ Appendix A Auxiliary Results
On the one hand, via (1.3), (B.11) and Lemma A.1(i), using that $(2+\delta)V_{0}^{\prime\prime}(\xi^{0}_{\star})\gamma_{1}>1+\frac{\delta}{2}$ if $\beta=1$ by (B.12),
@@
$S^{2,2}_{h}$ $\coloneqq\sum_{k=1}^{\lceil h^{-2}\rceil}\mathbb{E}\bigg[\Big(\frac{h^{-1}}{\lceil h^{-2}\rceil}\eta^{h}_{k}\Big)^{2}\bigg|\mathcal{F}^{h}_{k-1}\bigg]=\frac{h^{-2}}{\lceil h^{-2}\rceil}\frac{\operatorname{\mathbb{V}ar}\big((X_{h}-\xi^{h}_{\star})^{+}\big)}{(1-\alpha)^{2}},$
@@
$\Sigma_{n}=\frac{1}{\gamma_{n}}\sum_{k=1}^{n}\gamma_{k}^{2}\Pi_{k+1:n}^{2}.$ (B.18)
@@
$\lim_{\mathcal{H}\ni h\downarrow 0}S^{1,1}_{h}=\frac{\gamma_{1}\alpha}{1-\alpha}\lim_{n\uparrow\infty}\Sigma_{n},$ (B.19)
@@
$\frac{\gamma_{n}-\gamma_{n+1}}{\gamma_{n+1}}=\frac{\mathds{1}_{\beta=1}}{\gamma_{1}}\gamma_{n}+\operatorname{o}(\gamma_{n}),\quad\gamma_{n+1}-\gamma_{n}=\operatorname{o}(\gamma_{n})\quad\mbox{ and }\quad\gamma_{n}\gamma_{n+1}=\operatorname{o}(\gamma_{n}).$
@@
$\Sigma_{n+1}=\Sigma_{n}+\Big(1-\Big(2V_{0}^{\prime\prime}(\xi^{0}_{\star})-\frac{\mathds{1}_{\beta=1}}{\gamma_{1}}\Big)\Sigma_{n}\Big)\gamma_{n}+(\Sigma_{n}+1)\operatorname{o}(\gamma_{n}).$
@@
$\Sigma_{\star}\coloneqq\frac{1}{2V_{0}^{\prime\prime}(\xi^{0}_{\star})-\frac{\mathds{1}_{\beta=1}}{\gamma_{1}}}.$ (B.20)
@@
$\Delta\Sigma_{n}\coloneqq\Sigma_{n}-\Sigma_{\star},$ (B.21)
@@
$=\big(1-\mu\gamma_{n}+\operatorname{o}(\gamma_{n})\big)\Delta\Sigma_{n}+\operatorname{o}(\gamma_{n}),$
@@
$|\Delta\Sigma_{n+1}|\leq\big(1-(\mu+\varepsilon)\gamma_{n}\big)|\Delta\Sigma_{n}|+\varepsilon\gamma_{n}.$ (B.22)
@@
$\lim_{n\uparrow\infty}\Sigma_{n}=\Sigma_{\star}.$ (B.23)
Combining (B.19) and (B.23),
$S^{1,1}_{h}\mathrel{{\mathop{\longrightarrow}\limits^{\operatorname{\mathds{P}\text{-as}}}}}\frac{\alpha}{1-\alpha}\,\frac{\gamma_{1}}{2V_{0}^{\prime\prime}(\xi^{0}_{\star})-\frac{\mathds{1}_{\beta=1}}{\gamma_{1}}}\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
Given that $x\mapsto x^{+}$ is $1$-Lipschitz on $\mathbb{R}$,
@@
$\mathbb{E}[\operatorname{\mathbb{V}ar}(\varphi(Y,Z)|Y)]\leq\mathbb{E}[\operatorname{\mathbb{V}ar}(\varphi(Y,Z)|Y)]+\operatorname{\mathbb{V}ar}(\mathbb{E}[\varphi(Y,Z)|Y])=\operatorname{\mathbb{V}ar}(\varphi(Y,Z))<\infty.$
@@
$\mathbb{E}[(X_{h}-X_{0})^{2}]=h\,\mathbb{E}[\operatorname{\mathbb{V}ar}(\varphi(Y,Z)|Y)]\to 0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$ (B.24)
@@
$\operatorname{\mathbb{V}ar}\big((X_{h}-\xi^{h}_{\star})^{+}\big)\to\operatorname{\mathbb{V}ar}\big((X_{0}-\xi^{0}_{\star})^{+}\big)\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
$S^{2,2}_{h}\mathrel{{\mathop{\longrightarrow}\limits^{\operatorname{\mathds{P}\text{-as}}}}}\frac{\operatorname{\mathbb{V}ar}\big((X_{0}-\xi^{0}_{\star})^{+}\big)}{(1-\alpha)^{2}}\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
$\mathbb{E}\big[e^{h}_{k}\eta^{h}_{k}\big|\mathcal{F}^{h}_{k-1}\big]=\frac{\alpha}{(1-\alpha)^{2}}\,\mathbb{E}[(X_{h}-\xi^{h}_{\star})^{+}]\to\frac{\alpha}{(1-\alpha)^{2}}\,\mathbb{E}[(X_{0}-\xi^{0}_{\star})^{+}]\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$ (B.25)
@@
$\lim_{\mathcal{H}\ni h\downarrow 0}\frac{h^{-(1+\beta)}}{\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}\gamma_{k}\Pi_{k+1:\lceil h^{-2}\rceil}\mathbb{E}\big[e^{h}_{k}\eta^{h}_{k}\big|\mathcal{F}^{h}_{k-1}\big]=\frac{\alpha\,\mathbb{E}[(X_{0}-\xi^{0}_{\star})^{+}]\mathds{1}_{\beta=1}}{(1-\alpha)^{2}}\lim_{n\uparrow\infty}\widetilde{\Sigma}_{n},$ (B.26)
@@
$\lim_{n\uparrow\infty}{\sum_{k=1}^{n}\gamma_{k}\Pi_{k+1:n}}=\widetilde{\Sigma}_{\star}$. (B.27)
Finally, by (B.26) and (B.27),
$S^{1,2}_{h}=S^{2,1}_{h}\mathrel{{\mathop{\longrightarrow}\limits^{\operatorname{\mathds{P}\text{-as}}}}}\frac{\alpha\,\mathbb{E}[(X_{0}-\xi^{0}_{\star})^{+}]\mathds{1}_{\beta=1}}{V_{0}^{\prime\prime}(\xi^{0}_{\star})(1-\alpha)^{2}}\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
By the decomposition (B.1),
$\xi^{h}_{n}-\xi^{h}_{\star}=\frac{1}{V_{h}^{\prime\prime}(\xi^{h}_{\star})}a^{h}_{n}+\frac{1}{V_{h}^{\prime\prime}(\xi^{h}_{\star})}r^{h}_{n}+\frac{1}{V_{h}^{\prime\prime}(\xi^{h}_{\star})}\rho^{h}_{n}+\frac{1}{V_{h}^{\prime\prime}(\xi^{h}_{\star})}e^{h}_{n},$ (C.1)
recalling the definitions (B.3)–(B.5), with $\{a^{h}_{n},n\geq 1\}$ given by
$a^{h}_{n}=-\frac{1}{\gamma_{n}}\big(\xi^{h}_{n}-\xi^{h}_{\star}-(\xi^{h}_{n-1}-\xi^{h}_{\star})\big).$ (C.2)
From (2.1) and (C.1),
@@
$\frac{1}{n}\sum_{k=1}^{n}\frac{1}{\gamma_{k}}\big(\xi^{h}_{k}-\xi^{h}_{\star}-(\xi^{h}_{k-1}-\xi^{h}_{\star})\big)$ $=\frac{1}{n}\Big(\frac{1}{\gamma_{n}}(\xi^{h}_{n}-\xi^{h}_{\star})-\frac{1}{\gamma_{1}}(\xi^{h}_{0}-\xi^{h}_{\star})\Big)$ (C.3)
@@
$\mathbb{E}\bigg[\bigg|\frac{1}{n}\sum_{k=2}^{n}\Big(\frac{1}{\gamma_{k-1}}$ $-\frac{1}{\gamma_{k}}\Big)(\xi^{h}_{k-1}-\xi^{h}_{\star})\bigg|\bigg]$ (C.4)
@@
$\mathbb{E}\bigg[\bigg|\frac{1}{n}\sum_{k=1}^{n}\frac{1}{\gamma_{k}}a^{h}_{k}\bigg|\bigg]\leq\frac{C}{n\sqrt{\gamma_{n}}}.$ (C.5)
@@
$\frac{h^{-1}}{V_{h}^{\prime\prime}(\xi^{h}_{\star})\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}a^{h}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
$\frac{h^{-1}}{V_{h}^{\prime\prime}(\xi^{h}_{\star})\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}r^{h}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
Recalling that $\{\rho^{h}_{k},k\geq 1\}$ is a sequence of martingale increments, following (B.15),
@@
$\frac{h^{-1}}{V_{h}^{\prime\prime}(\xi^{h}_{\star})\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}\rho^{h}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
We start by checking the conditional Lindeberg condition. Note that Step 7.1 of the proof of Theorem 1.4, Appendix B, already guarantees that $\big\{\frac{h^{-1}}{\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}\eta^{h}_{k},h\in\mathcal{H}\big\}$ satisfies this condition. Moreover, by assumption, via (B.5) and (1.3),
@@
$\leq\frac{c_{\alpha}^{2+\delta}}{|V_{h}^{\prime\prime}(\xi^{h}_{\star})|^{2+\delta}\lceil h^{-2}\rceil^{\frac{\delta}{2}}}\to 0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0,$
@@
$S^{2,2}_{h}$ $\coloneqq\sum_{k=1}^{\lceil h^{-2}\rceil}\mathbb{E}\bigg[\Big(\frac{h^{-1}}{\lceil h^{-2}\rceil}\eta^{h}_{k}\Big)^{2}\bigg|\mathcal{F}^{h}_{k-1}\bigg]=\frac{h^{-2}}{\lceil h^{-2}\rceil}\frac{\operatorname{\mathbb{V}ar}\big((X_{h}-\xi^{h}_{\star})^{+}\big)}{(1-\alpha)^{2}},$
@@
$S^{1,1}_{h}\mathrel{{\mathop{\longrightarrow}\limits^{\operatorname{\mathds{P}\text{-as}}}}}\frac{\alpha(1-\alpha)}{f_{X_{0}}(\xi^{0}_{\star})^{2}}\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
Finally, using that $\lim_{\mathcal{H}\ni h\downarrow 0}V_{h}^{\prime\prime}(\xi^{h}_{\star})=V_{0}^{\prime\prime}(\xi^{0}_{\star})=(1-\alpha)^{-1}f_{X_{0}}(\xi^{0}_{\star})$ and (B.25), we obtain by Cesàro’s lemma that
$S^{1,2}_{h}=S^{2,1}_{h}\mathrel{{\mathop{\longrightarrow}\limits^{\operatorname{\mathds{P}\text{-as}}}}}\frac{\alpha}{1-\alpha}\frac{\mathbb{E}[(X_{0}-\xi^{0}_{\star})^{+}]}{f_{X_{0}}(\xi^{0}_{\star})}\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
$U_{\ell}\coloneqq\begin{pmatrix}\frac{1}{(KM^{\ell-1}(M-1))^{\frac{1}{2}}}\sum_{k=KM^{\ell-1}+1}^{KM^{\ell}}\Phi^{(k)}(Y)\\ \frac{1}{(KM^{\ell-1})^{\frac{1}{2}}}\sum_{k=1}^{KM^{\ell-1}}\Phi^{(k)}(Y)\end{pmatrix}\mathrel{{\mathop{\longrightarrow}\limits^{\mathcal{L}}}}\mathcal{N}\big(0,\mathbb{E}[\Phi^{(1)}(Y)^{2}|Y]I_{2}\big),\quad\mbox{ as }\quad\ell\uparrow\infty,$
@@
$h_{\ell}^{-\frac{1}{2}}(X_{h_{\ell}}-X_{h_{\ell-1}})\mathrel{{\mathop{\longrightarrow}\limits^{\mathcal{L}}}}\mathcal{N}\big(0,(M-1)\operatorname{\mathbb{V}ar}(\varphi(Y,Z)|Y)\big)\quad\mbox{ as }\quad\ell\uparrow\infty,$ (D.1)
@@
$\mathbb{E}[\operatorname{\mathbb{V}ar}(\varphi(Y,Z)|Y)]\leq\mathbb{E}[\operatorname{\mathbb{V}ar}(\varphi(Y,Z)|Y)]+\operatorname{\mathbb{V}ar}(\mathbb{E}[\varphi(Y,Z)|Y])=\operatorname{\mathbb{V}ar}(\varphi(Y,Z))<\infty,$ (D.2)
so that $\operatorname{\mathbb{V}ar}(\varphi(Y,Z)|Y)<\infty$ -$\operatorname{\mathds{P}\text{-as}}$. This concludes the proof.
@@
$\mathbb{P}(X_{h_{\ell-1}}\leq\xi<X_{h_{\ell}})$ $=\mathbb{P}\big(X_{h_{\ell-1}}\leq\xi<X_{h_{\ell-1}}+h_{\ell}^{\frac{1}{2}}G_{\ell}\big)$ (D.3)
@@
$\mathbb{P}(X_{h_{\ell}}\leq\xi<X_{h_{\ell-1}})=h_{\ell}^{\frac{1}{2}}\mathbb{E}[G_{\ell}^{-}f_{X_{h_{\ell-1}}|G_{\ell}}(\xi)]+h_{\ell}^{\frac{1}{2}}r_{\ell}^{-},$ (D.4)
@@
$\mathbb{E}[|G_{\ell}|f_{X_{h_{\ell-1}}|G_{\ell}}(\xi)]$ $=\mathbb{E}[|G_{\ell}|(f_{X_{h_{\ell-1}}|G_{\ell}}(\xi)-f_{G_{\ell}}(\xi))\mathds{1}_{|G_{\ell}|\leq C}]$ (D.5)
@@
Using that $\big\{g\mapsto f_{X_{h_{\ell-1}}|G_{\ell}=g}(\xi),\ell\geq 1\big\}$ are bounded uniformly in $\ell\geq 1$ and converge locally uniformly towards $g\mapsto f_{g}(\xi)$, by the dominated convergence theorem,